OpenStudy (anonymous):

what is the answer for this? 4x - 3y = 1 2x + 9y = 4

OpenStudy (anonymous):

multiply the first equation by 3 (4x - 3y = 1)3------>12x-9y=3 now add both equations 12x-9y=3 +2x+9y=4 ----------- 14x=7 x=7/14=1/2 now plug x value into one of the equation 2(1/2)+9y=4 1+9y=4 9y=4-1 y=3/9=1/3 so x=1/2 and y=1/3

OpenStudy (anonymous):

why 3?

OpenStudy (anonymous):

\[y=\frac{1}{3}\] not 3

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw :)

OpenStudy (anonymous):

\[a ^{2} x a ^{3} x a ^{-2}/ a ^{-3} x a ^{6}= how about this one please?

OpenStudy (anonymous):

\[\frac{a^{2+3-2}}{a^{-3+6}}\] \[\frac{a^3}{a^3}=1\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

24x ^{3} y ^{2} z ^{3}/ 4x ^{2} yz ^{2}= ?? help

OpenStudy (anonymous):

\[x ^{3} y ^{2} x ^{5} + 2x ^{7} y ^{-4} xy ^{6}\]

OpenStudy (anonymous):

dont post questions in the same thread